What is the value of log 1000 + 2log3 9-log5 125? Where 3 and 5 are the base perspectively.
Answer by oscargut(686)

log 1000 + 2log3 9-log5 125
log 1000 (the base is 10 because is ommited) so log 1000=3 because 10^3=1000
log3(9)=2 because 3^2=9
log5(125)=3 because 5^3=125
then log 1000 + 2log3 9-log5 125 = 3+2(2)-3=4
Answer:4